OpenStudy (anonymous):

How do you solve 5x + 12 = 6x - 8

OpenStudy (anonymous):

10x is your answer

OpenStudy (anonymous):

5x+12=6x-8 =5x-6x=12+8 =(-1x=20)-1 =x=-20